Solve for x. 4x(x-2)-5x(x-1)=2
Thank

Answer by rapaljer(4671)

4x(x-2)-5x(x-1)=2
4x^2 -8x -5x^2 + 5x =2
-x^2 -3x = 2

This is a quadratic equation, so you have to set the equation equal to zero. Because you have a NEGATIVE x^2 term on the left side, it will be a good thing to get everything on the RIGHT side by adding x^2 + 3x to each side:

-x^2 -3x+x^2 +3x= x^2 + 3x +2
0=x^2 + 3x +2

Factor the trinomial:
0=(x+2)(x+1)

Set each factor equal to zero:
x=-2 and x= -1.

For additional explanation on this type of problem, especially SOLVING QUADRTIC EQUATIONS (and maybe also FACTORING TRINOMIALS, please see my own website. Do a "Bing" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of the search list. On my Homepage near the top, look for the link "Basic, Intermediate and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in Chapter 2 for several topics on FACTORING and QUADRATIC EQUATIONS. In these sections, you will find my own non-traditional explanations, that my own students all said were a LOT easier to understand than traditional textbooks. See also the corresponding "MATH IN LIVING COLOR" pages in which problems are solved IN COLOR! I also have FREE videos on the website of me teaching my classes a few years ago. To access the videos, go back to my Homepage, and look for the link "Rapalje Videos in Living Color." Select "Basic Algebra", and select the video on "FACTORING." I don't have a video on Solving Quadratic Equations.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus